OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. (2 points) Question 1 options: 1) 203 2) 197 3) 168 4) 163

rishavraj (rishavraj):

\[a_{18} = a_1 + (n - 1) d\] plug values and get "d" i.e common difference

OpenStudy (anonymous):

i don"t understand this @rishavraj

rishavraj (rishavraj):

\[a_n = a_1 + (n - 1)d\] \[a_1 = -7\] \[a_{18} = a_1 + (n - 1)d\] so 95 = -7 + (17)d get "d" and then get 35th term by using the general equation mentiond above

OpenStudy (anonymous):

do u go to cnneitons 8gade

OpenStudy (anonymous):

@becca1